Prev Next

SW&IT Meeting Minutes 2021-09-06

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210906
LR20210823 - Bec, Matthieu 80.0
LR20210823 - Mechtley, Brandon 80.0
LR20210823 - Eychaner, Glenn 80.0
LR20210823 - Edmeades, David 80.0
LR20210823 - Grenz, Paul 10.0
LR20210823 - Hooper, Stephen 0.0
LR20210823 - Zhang, Yang 80.0
LR20210823 - Zhang, Xianyu 96.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • misc dms cruft removal. ks dms.mountain f34, mongo shuffle, hit some RAM issue for corner case mongo queries, bump VM spec and restarted - ok now
    • pb upgrade meeting logisitics with WR and others (3 time zones), exchanged notes with BM
    • IT8084 slice timestamp in web://info
    • discuss luci/x with GE/SH, created SSD 2022 placeholder https://wiki.lbto.org/Software/SummerShudown2022
    • IT expenditures follow up, looking okay, set a cal reminder
    • dms certificates expired (rpm out of date), set a cal reminder, how do we address others?
    • AO SX overcurrent 'analyis' - 2 lines of bash really - interesting findings though
  • This week
    • SWG reports, overdue
    • pb upgrade, catchup with Brandon, meet-up Aracteri et al. probably Thu or Fri around 12:00
    • luci activities catch-up with Glenn
    • "concern" MODS computer /restart and general snafu, plans to understand and address - particularly OSU

GCE

  • Last week:
    • IT 8454: LUCI problem with '/' drive becoming full
      • Fix is temporary; 25GB were freed by moving backups to lucix.luci and eliminating MariaDB binary logs. Permanent fix detailed in IT 8454.
      • M.Bec scheduling lucix.luci full-up test for SSD 2022.
    • Met with J.Power for overview of LUCI simulators and simulation
    • Sat in on LUCI2 warm test with D.Thompson
    • Last of the safety training
    • More learning of the IIF and Telemetry interfaces, needed to improve LUCI email alerting.
  • This week:
    • Sit in on LUCI2 on-telescope testing with D.Thompson. delayed by LUCI2 cold test fail
    • Prioritize and begin LUCI work
    • Give lucix.luci a once-over and compare it to luci.luci, just in case?
      • Check lucix bridged network to see if it can talk to the instrument. It can't take data, but it might be able to test instrument control.

YZ

  • Previous:
    • lbto-osco v0.8 release (plus reworked RPM build system, now installs systemd service unit on dms.mountain by default, don't have permission to rename the LBTO github repo's name!).
    • MODS IDL and software issues communicated with OSU again.
    • LBC bias value measure and adjustment with Alessandro and Olga.
    • TMS hardware configuration updates (in software), plus night time test with Heejoo (failed on first night, but subsequent night seemed to be a success).
    • Finished TCS refmem sampler (for tcs2) refactoring (working and tuning the WCS client now, PR to follow soon).
  • Next:
    • Finish DMS WCS refmem clients and create pull requests.
    • Wait on OSU's reply on several MODS issues (may set a wait limit on software and then move forward).
    • Start working on some of the TMS software issues and planned tasks.

DE

  • More MS licensing stuff, still nothing from CDW.
  • Systemd logrotate fun; will check on progress this week
  • Resource management for MB on vm1.mountain
  • Drive work with LD on mountain Synology
    • He checked warranty status/all out
    • He ordered more spares
This week:
  • Plan for upgrading hardware/os of backup host
    • Include more local storage for backups
    • More RAM?
  • Ready to set up purchase for new Syno gear once we get the go-ahead
  • Archive meeting
  • Begin offboarding ME
  • Dates for others?

BM

  • Last week:
    • Catch-up with AdSec overcurrent issues; AO group bi-weekly.
    • Catch-up with AdSec PB upgrade.
    • Met with JC re: AOS (pupil tests), got a very helpful half demonstration.
      • JC mentioned bandwidth issues when everything is running is normal (maybe due to the actuator sensor data for the visualization?), but the GUI was extraordinarily laggy Friday to the point of being inoperable.
  • This week:
    • Meet with MB (and Arcetri?) re: PB.
    • Continue AO sessions with Al.
    • Start looking at AOS, IDL procedures. Testing sandbox?
    • Follow-up with JC to understand AOS, quirks.

PSG

  • Last week:
    • Testing using project with scout software for Siemans D425 (test unit in lab).
    • Isuses with using existing project as the unit in the lab is vaerion 4.3, but project works with 4.1.
    • Unit on mountain checked, is version 4.1 so update should work.
  • This week:
    • Gustavo and Juan Pablo will go to the mountain and update the unit carfeully.
    • Generated project will be checked to ensure it is identical, only then will it be updated.
    • Testing then will be done if the update is successful.
Topic revision: r9 - 09 Sep 2021, PaulGrenz
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback